# Hermitian adjacency matrices with at most three distinct eigenvalues

### v1

@@ 2.2 Regular Oriented Graphs
Certainly $D$ must be a directed cycle. From Lemma 2.1 the eigenvalues are $1$ and $-2$, so $H=H_{\omega}(D)$ has minimal polynomial $x^{2}+x-2=0$. Thus we have
$H^{2}+H-2I=0.$
If $u,v$ are two vertices at distance 2 in $D$ then, since $D$ is a directed cycle, the $(u,v)$ entry of $H^{2}$ is in $\{\omega^{2},\omega^{4},\omega^{2}+\omega^{4}\}$ (the first two possibilities are generic and the third occurs exactly if $D$ is a 4-cycle). In any case, this implies that the $(u,v)$ entry of $H$ is non-zero, so $u$ and $v$ were adjacent, a contradiction. Hence, the only possibility is that $D$ is a directed 3-cycle, and indeed, this oriented graph has two $H_{\omega}$-eigenvalues. ∎
@@
Let $D$ is an $r$-regular oriented graph with exactly two distinct $H_{\omega}$-eigenvalues. If the underlying graph of $D$ is triangle-free, then $r=2$.
@@
From Lemma 2.1 we have $s=-2$. The formula for $\operatorname{tr}(D)$ gives
00 0 $=\operatorname{tr}(D)=rm-2(n-m).$
@@
00 0 $=\operatorname{tr}(D^{3})=r^{3}m-8(n-m).$
Since $n-m\neq 0$, substituting $rm=2(n-m)$ into the second equation gives $r^{2}=4$, hence $r=2$. ∎
@@
If $G$ is a connected $4$-regular triangle-free undirected graph and each pair of vertices at distance $2$ have three common neighbors, then $G$ is isomorphic to $K_{5,5}$ minus a perfect matching.
@@
The sets $C(w,x)$ and $C(w,y)$ cannot be disjoint because $w$ has degree 4. If $C(w,x)=C(w,y)$, then since $w$ has degree 4, we must also have $C(w,z)=C(w,x)$, and $|S|=2$, but then either vertex of $C(w,x)$ has four common neighbours with $v$, a contradiction.
So each of the six considered subsets of $S$ intersect in one vertex, and each of these vertices belongs to at most 3 of the subsets. This happens only if $|S|=4$ and each element of $S$ is adjacent to a distinct set of three vertices from $\{w,x,y,z\}$. Finally, each element of $S$ has one further neighbour, but each pair of vertices in $S$ has one further common neighbour, hence $S$ is the neighbourhood of a single vertex $v^{\prime}$, and $G$ is isomorphic to $K_{5,5}-M$. ∎
@@
$(3)\Rightarrow(1)$ and $(3)\Rightarrow(2)$ are trivial. $(2)\Rightarrow(1)$ is Lemma 2.7. If $(2)$ holds, then so does $(1)$, hence $r=-s=2$ and so the assumptions of Lemma 2.3 hold. Thus the assumptions of Lemma 2.8 hold, which implies $(3)$, so we have $(2)\Rightarrow(3)$.
@@
This is a triangle-free, regular oriented graph. Moreover, the underlying graph of $\widetilde{D}$ is connected since the underlying graph of $D$ is connected and not bipartite (see Theorem 3.4 from [1]). For any eigenvector $v$ of $H_{\omega}(D)$ with eigenvalue $\theta$, the vectors $(v,v)$ and $(v,-v)$ are eigenvectors of $H_{\omega}(\widetilde{D})$ with eigenvalues $\theta$ and $-\theta$ respectively. The set
$\{(v,v)\cup(v,-v):v\mbox{ is an eigenvector of }H_{\omega}(D)\}$
spans the eigenvectors of $H_{\omega}(\widetilde{D})$, hence $H_{\omega}(\widetilde{D})$ has exactly two distinct eigenvalues, $2$ and $-2$. Now, since $(2)\Rightarrow(3)$ and since $(2)$ holds for $\widetilde{D}$, the underlying graph of $\widetilde{D}$ must be isomorphic to $K_{5,5}-M$. This in turn implies that the underlying graph of $D$ is isomorphic to $K_{5}$. Up to isomorphism there is only a single regular tournament of order 5. We computed its $H_{\omega}$-eigenvalues directly and found that they are all distinct, a contradiction. ∎
@@ 2.3 Regular Oriented Graphs with $r\geq 3$
The remaining candidates for oriented graphs with two distinct $H_{\omega}$-eigenvalues are $r$-regular and have eigenvalues $r$ and $-2$. We show that such oriented graphs do not exist. This will complete the proof of Theorem 1.1.
@@
Let $D$ be an $r$-regular oriented graph with two distinct $H_{\omega}$-eigenvalues $r$ and $s$, where $r>s$. If $s=-2$, then $r\leq 2$.
@@
Let $H=H_{\omega}(D)$. Since $s=-2$, the minimal polynomial of $H$ is $(x-r)(x+2)$ and therefore we have
@@
Since $\omega^{3}=-1$ the right-hand side is real, so the left-hand side is zero. We conclude that $a=b$ and $a-c=r-2.$ Since $c$ is non-negative, $a\geq r-2$, so there are at least $r-2$ walks of length $2$ with value $1$ from $u$ to $v$. Since $b=a$ there are also at least $r-2$ walks of length $2$ with value $\omega^{2}$ from $u$ to $v$.
To show that $r\leq 4$ we argue that there are at most two $2$-walks of value $1$ from $u$ to $v$. The orientations of $2$-walks from $u$ to $v$ with value $1$ are $u\rightarrow x\leftarrow v$ and $u\leftarrow x\rightarrow v$, which we will call absorbing and repelling, respectively. Since the only $2$-walk with value $\omega^{2}$ is the directed $2$-walk, there must be $r-2$ directed $2$-walks from $u$ to $v$. In particular, these paths contribute $r-2$ incoming edges to $v$, and combined with the arc $(u,v)$, this means $v$ can only have one further incoming arc. Thus, there must be at most one repelling $2$-walk from $u$ to $v$. Similarly, the $r-2$ directed 2-walks contribute $r-2$ outgoing edges from $u$, and combined with the arc $(u,v)$, this means there must be at most one absorbing 2-walk from $u$ to $v$.
@@ 3 Three Distinct Eigenvalues
In counterpoint to the previous section, we show that oriented graphs with three distinct $H_{\omega}$-eigenvalues are plentiful. The results of this section are a generalization of results from [2].
@@
where $J$ is the all ones matrix and $B-I$ is a skew-symmetric $(1,-1)$-matrix.
A Hadamard matrix is a square $(1,-1)$-matrix of order $n$ such that $HH^{T}=nI$. A skew-Hadamard matrix is a Hadamard matrix for which $H-I$ is skew-symmetric. Many constructions of skew-Hadamard matrices are known, see [9] for a survey of constructions. For us, the important fact is that there are infinitely many integers $k$ for which a skew-Hadamard matrix of order $4k$ exists. A regular tournament is a regular oriented complete graph.
@@
For each positive integer $k$ there is a regular tournament $T$ of order $14k-$ 1 for which $H_{\omega}(T)$ has exactly three distinct eigenvalues.
@@
Let $A$ be a skew-Hadamard matrix of order $n=4k$. Without loss of generality we may assume that the first row of $A$ is $j^{\intercal}$, the all one’s vector. Now, remove the first row and the first column of $A$ to obtain a matrix $B$ so that $B-I$ is a skew-symmetric matrix of order $14k-$ 1 . Note that in every row and column of $B-I$ there are $12k-$ 1, entries 1 and $12k-$ 1, entries $-1$. Clearly, $BB^{\intercal}=nI-J$. Thus $B^{2}=J-nI$. This yields that $BJ=JB$. Since $B-I$ is skew-symmetric, we have $(B-I)^{\intercal}=I-B$ and so we find that $B+B^{\intercal}=2I$.
Now let $T$ be the $(2k-1)$-regular tournament whose Hermitian adjacency matrix is
@@
Note that $T$ has order $14k-$ 1. We have
@@
It follows that the all one’s vector of length $n-1$ is an eigenvector of $B-I$ with eigenvalue 0, and the remaining eigenvectors are orthogonal to $J$, hence the remaining eigenvalues of $B-I$ are ±$\pm i\sqrt{n-1}$. Since $\operatorname{tr}(B-I)=0$, this latter pair of eigenvalues have equal multiplicity $\frac{n-2}{2}$.
Since $BJ=JB$, the three matrices $B-I,J$ and $I$ are simultaneously diagonalizable (note that $B$ is diagonalizable because $iB$ is Hermitian). Thus the eigenvalues of $H$ are $\frac{n-2}{2}$ of multiplicity 1 and
$-\frac{1}{2}\pm\frac{\sqrt{3(n-1)}}{2},$
each with multiplicity $\frac{n-2}{2}$. Hence $T$ is of order $n-1$ and has 3 distinct eigenvalues and the proof is complete. ∎
@@
Let $T$ be a regular tournament of order $n-1$ with 3 distinct $H_{\omega}$-eigenvalues. Then there is a skew-Hadamard matrix of order $n$.
@@
where $B-I$ is a skew-symmetric $(1,-1)$-matrix, whose all diagonal elements are zero. Since $T$ is a regular tournament of order $n-1$, one of the eigenvalues of $T$ is $\frac{n-2}{2}$. If $\theta$ is an eigenvalue of $B-I$, then $-\theta$ is an also an eigenvalue of $B-I$. Thus the three distinct eigenvalues of $H$ are $\frac{n-2}{2}$, with multiplicity 1, and $-\frac{1}{2}\pm\frac{\sqrt{3}}{2}\theta$, each with multiplicity $\frac{n-2}{2}$. Since $\operatorname{tr}(H^{2})$ is twice the number of edges we have
$\Big{(}\frac{n-2}{2}\Big{)}^{2}+\frac{n-2}{2}\Big{(}-\frac{1}{2}+\frac{\sqrt{3}}{2}\theta\Big{)}^{2}+\frac{n-2}{2}\Big{(}-\frac{1}{2}-\frac{\sqrt{3}}{2}\theta\Big{)}^{2}=(n-1)(n-2),$
hence we find that $\theta=\sqrt{n-1}$.
Since $i(B-I)$ is a Hermitian matrix, it is diagonalizable and so $B-I$ is diagonalizable. Thus the minimal polynomial of $B-I$ is $x(x-\sqrt{n-1}\,i)(x+\sqrt{n-1}\,i)$. This implies that
@@
Zero is an eigenvalue of multiplicity 1 for the matrix $B-I$ and $j$ is its corresponding eigenvector. Moreover $(B-I)^{2}+(n-1)I$ is a symmetric matrix such that its every column is an eigenvector of zero for the matrix $B-I$, so we conclude that $(B-I)^{2}+(n-1)I=cJ$, for some suitable $c$.
Since $(B-I)j=0$, we conclude that $(n-1)j=cJj=c(n-1)j$. Therefore $c=1$ and $(B-I)^{2}=J-(n-1)I$. Since $B-I$ is a skew-symmetric matrix, we have
@@
This equation implies that the matrix $A$ obtained from $B$ by adding a column of $-1$’s and then a row of $1$’s is a skew-Hadamard matrix of order $n$. ∎

### v2

@@ 2.2 Regular Oriented Graphs with $r\leq 2$
Certainly $D$ must be a directed cycle. From Lemma 2.1 the eigenvalues are $1$ and 2-2. If $D$ is not a triangle then $D$ contains an induced path on three vertices. Regardless of the orientation, $\sqrt{2}$ is an $H_{\omega}$-eigenvalue of this induced path, contradicting interlacing. ∎
@@
Let $D$ be an $r$-regular oriented graph with exactly two distinct $H_{\omega}$-eigenvalues. If the underlying graph of $D$ is triangle-free, then $r=2$.
@@
From Lemma 2.1 we have s=2s=-2. The formula for $\operatorname{tr}(D)$ gives
$0$ $=\operatorname{tr}(D)=rm-2(n-m).$
@@
$0$ $=\operatorname{tr}(D^{3})=r^{3}m-8(n-m).$
Since $n-m\neq 0$, substituting rm=2(nm)rm=2(n-m) into the second equation gives $r^{2}=4$, hence $r=2$. ∎
We let $K_{5,5}-M$ denote the graph obtained from $K_{5,5}$ by deleting the edges of a perfect matching.
@@
If $G$ is a connected $4$-regular triangle-free undirected graph and each pair of vertices at distance $2$ have three common neighbors, then $G$ is isomorphic to $K_{5,5}-M.$
@@
The sets $C(w,x)$ and $C(w,y)$ cannot be disjoint because $w$ has degree 4. If $C(w,x)=C(w,y)$, then since $w$ has degree 4, we must also have $C(w,z)\cap C(w,x)\neq\emptyset$, but then a vertex in $C(w,z)\cap C(w,x)$ has four common neighbours with $v$, a contradiction.
So each pair of the six considered subsets of $S$ intersect in one vertex, and each of these vertices belongs to at most 3 of the subsets. This happens only if $|S|=4$ and each element of $S$ is adjacent to a distinct set of three vertices from $\{w,x,y,z\}$. Finally, each element of $S$ has one further neighbour, but each pair of vertices in $S$ has one further common neighbour, hence $S$ is the neighbourhood of a single vertex $v^{\prime}$, and $G$ is isomorphic to $K_{5,5}-M$. ∎
@@
$(3)\Rightarrow(1)$ and $(3)\Rightarrow(2)$ are trivial. $(2)\Rightarrow(1)$ is Lemma 2.7. If $(2)$ holds, then so does $(1)$, hence r=s=2r=-s=2 and so the assumptions of Lemma 2.3 hold. Thus the assumptions of Lemma 2.8 hold, which implies $(3)$, so we have $(2)\Rightarrow(3)$.
@@
[…]
$\{(v,v),(v,-v):v\mbox{ is an eigenvector of }H_{\omega}(D)\}$
[…]
@@ 2.3 Regular Oriented Graphs with $r\geq 3$
The remaining candidates for oriented graphs with two distinct $H_{\omega}$-eigenvalues are $r$-regular and have eigenvalues $r$ and 2-2. We show that such oriented graphs do not exist. This will complete the proof of Theorem 1.1.
@@
Let $D$ be an $r$-regular oriented graph with two distinct $H_{\omega}$-eigenvalues $r$ and $s$, where $r>s$. If s=2s=-2, then $r\leq 2$.
@@
Let $H=H_{\omega}(D)$. Since s=2s=-2, the minimal polynomial of $H$ is (xr)(x+2)(x-r)(x+2) and therefore we have
@@
Since $\omega^{3}=-1$ the right-hand side is real, so the left-hand side is zero. We conclude that $a=b$ and ac=r2.a-c=r-2. Since $c$ is non-negative, $a\geq r-2$, so there are at least r2r-2 walks of length $2$ with value $1$ from $u$ to $v$. Since $b=a$ there are also at least r2r-2 walks of length $2$ with value $\omega^{2}$ from $u$ to $v$.
To show that $r\leq 4$ we argue that there are at most two $2$-walks of value $1$ from $u$ to $v$. The orientations of $2$-walks from $u$ to $v$ with value $1$ are $u\rightarrow x\leftarrow v$ and $u\leftarrow x\rightarrow v$, which we will call absorbing and repelling, respectively. Since the only $2$-walk with value $\omega^{2}$ is the directed $2$-walk, there must be r2r-2 directed $2$-walks from $u$ to $v$. In particular, these paths contribute r2r-2 incoming edges to $v$, and combined with the arc $(u,v)$, this means $v$ can only have one further incoming arc. Thus, there must be at most one repelling $2$-walk from $u$ to $v$. Similarly, the r2r-2 directed 2-walks contribute r2r-2 outgoing edges from $u$, and combined with the arc $(u,v)$, this means there must be at most one absorbing 2-walk from $u$ to $v$.
@@
2.4 Extension to Mixed Graphs
In this section we sketch the extension of Theorem 1.1 to mixed graphs. The arguments are a simple extension of the above and yield only one additional non-trivial example.
Recall that a mixed graph $D=(V,A,E)$ consists of a vertex set $V$ a set $A$ of arcs and a set $E$ of edges. As described in the introduction we extend the definition of $H_{\omega}(D)$ so that $H_{\omega}(D)_{u,v}=1$ when $\{u,v\}$ is an edge of $D$.
Theorem 2.12.
If $D$ is a connected mixed graph and $H_{\omega}(D)$ has exactly two distinct eigenvalues, then $D$ is one of the following.
  1. (a)
    An complete graph on $n\geq 2$ vertices cospectral to the unoriented complete graph.
  2. (b)
    A 4-cycle containing a directed path of length three and one edge.
  3. (c)
    One of the four oriented graphs from Theorem 1.1.
Proof.
As before we let $H=H_{\omega}(D)$ and let $s<0<r$ be the two distinct eigenvalues of $H$.
Note that Lemma 2.2 holds for mixed graphs. The only difference in the proof is when showing that $(r+s)$ must be an integer. There we must consider five types of walks of length 2 between $u$ and $v$ instead of three, but the same technique applies.
Next, note that the bound in Lemma 2.1 still holds and is sharp if $D$ has any undirected edges. So if $D$ is a mixed graph we have $s\in\{-1,-\sqrt{2},-\sqrt{3}\}$.
Suppose s=1s=-1. Interlacing implies that each three vertex induced subgraph of $D$ is a triangle with at least one edge and no directed path of length 2. It follows that we are in case $(a)$.
Suppose $s=-\sqrt{2}$. The graph underlying $D$ is cycle. Lemma 2.3 no longer holds but the argument in its proof implies that any pair of vertices at distance 2 in $D$ must have a number of common neighbours equal to $2a+3b$ for non-negative integers $a$ and $b$. So $C_{4}$ is the only new possibility for the underlying graph. The mixed orientation stated in case $(b)$ is easily verified to be the unique choice with two eigenvalues.
Suppose $s=-\sqrt{3}$. Note first that each mixed orientation of the triangle has an eigenvalue outside the interval $\sqrt{3},\sqrt{3}]$, so interlacing implies the graph underlying $D$ is triangle-free. Now in the underlying graph all vertices have degree 3 and each pair of vertices at distance 2 must have two or three common neighbours.
If there is a pair of vertices with three common neighbours then the argument used in the proof of Theorem 2.5 shows that $D$ must be the oriented $K_{3,3}$ found there. So we assume all vertices of $D$ have two common neighbors.
It is easy to verify that the cube is the only connected cubic triangle-free graph in which each pair of vertices at distance 2 have two common neighbors.
Suppose we have a mixed orientation of the cube with two $H_{\omega}$-eigenvalues. Each of the six 4-cycles must be the mixed orientation of $C_{4}$ mentioned above which contains one undirected edge. So this mixed cube has exactly three undirected edges and they form a maximal matching. The two remaining vertices uncovered by this matching are each required to be the center of three directed paths of length 2, one for each incident 4-cycle, but this is impossible. ∎
@@ 3 Three Distinct Eigenvalues
In counterpoint to the previous section, we show that oriented graphs with three distinct $H_{\omega}$-eigenvalues are plentiful. The results of this section are a generalization of results from [dCGKPM92].
@@
where $J$ is the all ones matrix and BIB-I is a skew-symmetric (1,1)(1,-1)-matrix.
A Hadamard matrix is a square (1,1)(1,-1)-matrix of order $n$ such that $HH^{T}=nI$. A skew-Hadamard matrix is a Hadamard matrix for which HIH-I is skew-symmetric. Many constructions of skew-Hadamard matrices are known, see [KS08] for a survey of constructions. For us, the important fact is that there are infinitely many integers $k$ for which a skew-Hadamard matrix of order $4k$ exists. A regular tournament is a regular oriented complete graph.
@@
For each positive integer $k$ there is a regular tournament $T$ of order 4k14k-1 for which $H_{\omega}(T)$ has exactly three distinct eigenvalues.
@@
Let $A$ be a skew-Hadamard matrix of order $n=4k$. Without loss of generality we may assume that the first row of $A$ is $j^{\intercal}$, the all one’s vector. Now, remove the first row and the first column of $A$ to obtain a matrix $B$ so that BIB-I is a skew-symmetric matrix of order 4k1.4k-1. Note that in every row and column of BIB-I there are 2k12k-1 entries 1 and 2k12k-1 entries 1-1. Clearly, $BB^{\intercal}=nI-J$ and $BJ=JB$. Since BIB-I is skew-symmetric, we have $(B-I)^{\intercal}=I-B$ and so we find that $B+B^{\intercal}=2I$.
Now let $T$ be the (2k1)(2k-1)-regular tournament whose Hermitian adjacency matrix is
@@
Note that $T$ has order 4k14k-1. We have
@@
It follows that the all one’s vector of length n1n-1 is an eigenvector of BIB-I with eigenvalue 0, and the remaining eigenvectors are orthogonal to $J$, hence the remaining eigenvalues of BIB-I are $\pm i\sqrt{n-1}$. Since $\operatorname{tr}(B-I)=0$, this latter pair of eigenvalues have equal multiplicity $\frac{n-2}{2}$.
Since $BJ=JB$, the three matrices BI,JB-I,J and $I$ are simultaneously diagonalizable (note that $B$ is diagonalizable because $iB$ is Hermitian). Thus the eigenvalues of $H$ are $\frac{n-2}{2}$ of multiplicity 1 and
$\frac{1}{2}\pm\frac{\sqrt{3(n-1)}}{2},$
each with multiplicity $\frac{n-2}{2}$. Hence $T$ is of order n1n-1 and has 3 distinct eigenvalues and the proof is complete. ∎
@@
Let $T$ be a regular tournament of order n1n-1 with $3$ distinct $H_{\omega}$-eigenvalues. Then there is a skew-Hadamard matrix of order $n$.
@@
where BIB-I is a skew-symmetric (1,1)(1,-1)-matrix, whose all diagonal elements are zero. Since $T$ is a regular tournament of order n1n-1, one of the eigenvalues of $T$ is $\frac{n-2}{2}$. If $\theta$ is an eigenvalue of BIB-I, then $\theta$ is also an eigenvalue of BIB-I. Thus the three distinct eigenvalues of $H$ are $\frac{n-2}{2}$, with multiplicity 1, and $\frac{1}{2}\pm i\frac{\sqrt{3}}{2}\theta$, each with multiplicity $\frac{n-2}{2}$. Since $\operatorname{tr}(H^{2})$ is twice the number of edges we have
$\Big(\frac{n-2}{2}\Big)^{2}+\frac{n-2}{2}\Big(-\frac{1}{2}+i\frac{\sqrt{3}}{2}\theta\Big)^{2}+\frac{n-2}{2}\Big(-\frac{1}{2}-i\frac{\sqrt{3}}{2}\theta\Big)^{2}=(n-1)(n-2),$
hence we find that $\theta=i\sqrt{n-1}$.
Since i(BI)i(B-I) is a Hermitian matrix, it is diagonalizable and so BIB-I is diagonalizable. Thus the minimal polynomial of BIB-I is $x(x-\sqrt{n-1}\,i)(x+\sqrt{n-1}\,i)$. This implies that
@@
Zero is an eigenvalue of multiplicity 1 for the matrix BIB-I and $j$ is its corresponding eigenvector. Moreover $(B-I)^{2}+(n-1)I$ is a symmetric matrix such that its every column is an eigenvector of zero for the matrix BIB-I, so we conclude that $(B-I)^{2}+(n-1)I=cJ$, for some suitable $c$.
Since (BI)j=0(B-I)j=0, we conclude that (n1)j=cJj=c(n1)j(n-1)j=cJj=c(n-1)j. Therefore $c=1$ and $(B-I)^{2}=J-(n-1)I$. Since BIB-I is a skew-symmetric matrix, we have
@@
This equation implies that the matrix $A$ obtained from $B$ by adding a column of 1-1’s and then a row of $1$’s is a skew-Hadamard matrix of order $n$. ∎